OpenStudy (anonymous):

find the volume below the surface z=x^2+y^2 above z=0 and inside the cylinder x^2+y^2=2y (ANSWER=3π/2)

OpenStudy (p0sitr0n):

switch to cylindrical coordinates

OpenStudy (p0sitr0n):

integral from 0 -> 2PI integral from 0-> root2 x^2+y^2 \[\int\limits_{0}^{2\pi}\int\limits_{0}^{\sqrt{2}}zrdrd \theta \]

OpenStudy (anonymous):

be carefull, paraboloid and cilinder are not concentrical

ganeshie8 (ganeshie8):

\[\iiint 1 dV\]

OpenStudy (anonymous):

why integral 0 to 2π

ganeshie8 (ganeshie8):

Oh thats a good point @myko center of shadow in xy plane gets shifted to (0, 1)

ganeshie8 (ganeshie8):

x^2+y^2=2y x^2 + (y-1)^2 = 1